1.    What is the angular acceleration of a ball that starts at rest and increases its angular velocity uniformly to 5 rad/s in 10 s? 



1
Expert's answer
2021-12-21T12:16:15-0500
"\omega=\omega_0+\alpha t,""\alpha=\dfrac{\omega-\omega_0}{t}=\dfrac{5\ \dfrac{rad}{s}-0}{10\ s}=0.5\ \dfrac{rad}{s^2}."